# IDEA: Enhancing the Rule Learning Ability of Large Language Model Agent through Induction, Deduction, and Abduction

### v4

@@ Abstract
While large language models (LLMs) have been thoroughly evaluated for deductive and inductive reasoning, their proficiency in abductive reasoning and holistic rule learning in interactive environments remains less explored. We introduce RULEARN, a novel benchmark specifically designed to assess the rule-learning abilities of LLM agents in interactive settings. In RULEARN, agents strategically interact with simulated environments to gather observations, discern patterns, and solve complex problems. To enhance the rule-learning capabilities for LLM agents, we propose IDEA, a novel reasoning framework that integrates the process of Induction, DEduction, and Abduction. The IDEA agent generates initial hypotheses from limited observations through abduction, devises plans to validate these hypotheses or leverages them to solve problems via deduction, and refines previous hypotheses using patterns identified from new observations through induction, dynamically establishing and applying rules that mimic human rule-learning behaviors. Our evaluation of the IDEA framework, which involves five representative LLMs, demonstrates significant improvements over the baseline. Furthermore, within this framework, our comparison with 50 human participants reveals notable discrepancies in rule-learning behaviors. LLM agents tend to generate plausible initial hypotheses but struggle to refine them through interaction. Conversely, humans, despite sometimes overlooking initial details, excel at incorporating feedback and continuously improving their hypotheses. We believe our benchmark, RULEARN, will serve as a valuable and challenging resource, and that the IDEA framework will provide crucial insights for the development of LLM agents capable of human-like rule learning in real-world scenarios. We will release our code and data upon acceptance of the paper.
@@ 1 Introduction
One major pillar of human intelligence is the ability to discern rules and apply them. We identify patterns, formulate hypotheses, and refine them by interacting with the environment. This exploratory process traditionally involves three stages: abduction, deduction, and induction. According to Charles Peirce’s definition (Frankfurt, 1958; Peirce, 1974), the rule-learning loop typically begins with an explanatory hypothesis that arises from abduction. This is followed by iterative experiments guided by the hypothesis, known as deduction, which leads to the modification and refinement of the hypothesis through induction (As shown in Figure 1). Whether the hypothesis emerges from rigorous logical reasoning or a more creative process is debated, but the cyclic nature of this reasoning is pivotal in advancing our understanding and problem-solving capabilities in various disciplines. This establishes the foundation of how humans learn the rules of the world and apply them to effect change.
For such cyclic reasoning to function effectively, it requires an interactive foundation. In the real world, the observations we reason from are gathered through interaction. Then, through alternating cycles of refining hypothesis and interaction, we iteratively collect evidence and refine our hypotheses, drawing ever closer to the truth. The dynamic interplay between interaction and different stages of the reasoning loop within an interactive environment is the key to humans’ reasoning effectiveness. However, recent studies focusing on reasoning often lack this dynamic and interdependent nature, as abduction, induction, and deduction are tested one at a time in non-interactive environments (Bowen et al., 2024; Wang et al., 2023; Saparov et al., 2024; Liu et al., 2024).
To address this gap, we introduce RULEARN, a benchmark designed to evaluate the rule-learning abilities of LLM agents in interactive environments. RULEARN consists of puzzles with hidden rules set in a text-based environment, where agents begin exploration without prior knowledge of these rules. The agent’s objective is to solve the puzzle by interacting with the environment. At each step, the agent selects an action from a predefined set, executes it, and gathers observations to gain insights into the environment. Successfully solving the puzzles requires the agent to strategically select actions, efficiently gather pattern-revealing observations, and accurately reason from them to infer the hidden rules. RULEARN presents substantial challenges, as agents must rely entirely on observations generated by their chosen actions to discern rules. If these observations fail to reveal a clear pattern, the agent is likely to fail.
@@
Figure 2: A simplified puzzle in the RULEARN benchmark and the IDEA agent’s workflow (in real puzzles, agents have fewer initial observations and more complex rules). The agent generates an initial hypothesis through abduction, develops an exploration plan via deduction, and refines its hypothesis using induction. For example, the IDEA agent first hypothesizes that the password is the number of the blue paintings, tests this by entering code 003, and adjusts its strategy based on the results.
@@
We design three types of environments within RULEARN to evaluate the rule-learning ability in different scenarios: (1) The Function Operator: The agent is tasked with determining the coefficients of mathematical functions defined by hidden expressions. Agents can assign various values to the input variables and observe the outputs, using this information to hypothesize the function’s form. The challenge lies in efficiently selecting input values that reveal the underlying structure and accurately computing the coefficients based on limited observations. (2) The Escape Room: The agent is tasked with deciphering the passcode to exit an escape room. A hidden rule determines how the objects in the room infer the passcode digits. Agents interact with these objects to gather clues and input passcodes into the door. The door provides hints indicating which digits are correct. Based on these hints, agents formulate hypotheses and infer the relationship between the objects and the passcode. (3) The Reactor: The agent is tasked with synthesizing target strings using a reactor with a hidden string-combining rule. Agents need to experiment with different inputs and analyze outputs to deduce the reactor’s transformation rule and achieve the desired outcome. Detailed puzzle definitions are provided in § 3.
@@
To tackle the challenge in RULEARN, we introduce IDEA, a novel reasoning framework that integrates the process of Induction, DEduction, and Abduction. The IDEA agent employs these reasoning processes iteratively to explore the environments, learn rules, and achieve goals. In the abduction phase, the IDEA agent generates an initial hypothesis from limited observations. During the deduction phase, the IDEA agent creates and executes plans to meet objectives or test its hypothesis. In the induction phase, the IDEA agent refines its hypothesis based on new observations, enhancing their accuracy and robustness. This iterative cycle enables the LLM agent to continually improve the learned rules through environmental feedback. An overview of how the IDEA agent solves puzzles in RULEARN is shown in Figure 2.
We evaluat the IDEA framework using five popular LLMs—GPT-3.5-Turbo, GPT-4o, Gemma-7B, Llama3-8B and Llama3-70B. The experimental results demonstrate that IDEA significantly improves rule-learning proficiency and puzzle-solving performance of the LLM agent within the RULEARN benchmark. The IDEA agent increases success rates by approximately 10% compared to the baseline LLM agent. Specifically, the action of the IDEA agent aims to validate or leverage current hypothesis, providing important guidance in an environment with hidden rules. Without the guidance of the hypothesis, the LLM agent will always choose the most direct actions attempting to quickly solve the puzzle. Such actions, taken without understanding the hidden rules, are akin to navigating without a compass—efforts are made, but they often lead to ineffective or misguided outcomes. Based on our experiments, the IDEA agent achieves 47.6% fewer repeated actions, yielding more diverse observations and finally solving the puzzles with a better understanding of the hidden rule. These improvements suggest that IDEA enables LLMs to better adapt and learn rules in unfamiliar environments, narrowing the gap between human and AI rule-learning abilities.
Within the IDEA framework, we conduct a study to compare the rule-learning behaviors and performances between LLMs and 50 human participants solving puzzles. Our findings reveal that, despite the improvements brought by the IDEA framework, LLMs still face the following challenges when compared to human participants: (1) LLM agents continue to struggle with efficiently exploring unfamiliar environments. The new observations gathered usually do not provide enough information to help reveal the underlying rules, moreover, LLM agents tend to repeat previous actions that do not yield new information. (2) LLMs fall short in deducting valid plans to verify current hypotheses and guide future exploration. (3) LLMs fall short to correct previous hypotheses when they contradict new observations and are less capable of refining a hypothesis to make it more robust and encompassing all observations. These findings provide valuable insights for the development of LLM agents capable of human-like rule learning in real-world scenarios.
@@ 2 Related Works
2.1 LLM agent
Agents driven by large language models equipped with memory, tool use, planning, and reasoning capabilities have shown significant advancements in understanding complex tasks and achieving notable performance (Wang et al., 2024b). Techniques like memory summarization (Chen et al., 2023; Zhou et al., 2023; Wang et al., 2024a) and retrieval (Andreas, 2022; Park et al., 2023; Zhong et al., 2023) assist LLM agents in filtering relevant information based on the current context (Zhang et al., 2024). Additionally, the integration of tool use enables LLM agents to perform better in knowledge-grounded tasks. For instance, some studies (Nakano et al., 2022; Lu et al., 2023; Shi et al., 2023a) have enabled LLM agents to access up-to-date information from external databases to provide correct answers, while others (Schick et al., 2023) have facilitated the use of various tools like calculators to enhance solution accuracy. Planning modules (Yuan et al., 2023; Shen et al., 2023) allow these agents to decompose complex tasks into simpler components, achieving significant improvements. Moreover, employing search algorithms (Yao et al., 2023; Besta et al., 2024) helps LLM agents reach solutions in fewer steps. Building on this foundation, we are interested in whether LLM agents can derive new knowledge–learning rules through interacting with the environment, rather than merely answering questions. Recent studies have begun investigating different types of reasoning in LLMs, including abduction, deduction, and induction (Bowen et al., 2024; Wang et al., 2023; Saparov et al., 2024; Cheng et al., 2024; Yang et al., 2024). However, existing research mostly focuses on isolated aspects of the rule-learning process. Consequently, the comprehensive rule-learning abilities of LLMs in interactive environments remain largely underexplored, a topic that will be further examined in Section 2.2.
2.2 Rule Learning
Rule learning—involving evidence gathering, pattern identification, hypothesis formulation, and validation—is gaining focus in recent research. Yang et al. (2023) and Zhu et al. (2024) improved question-answering systems by formulating optimal response rules, while Shi et al. (2023b) advanced event prediction by enabling LLMs to hypothesize causes of previous events. Some works also evaluate LLMs rule-learning abilities (Liu et al., 2024). However, existing studies adopt static, non-interactive settings for rule learning, where LLMs are provided with complete information to formulate rules. This approach has two significant limitations: it contrasts with real-world rule learning, which involves actively and strategically gathering evidence, and it restricts agents from conducting experiments to validate the rules they learn. In real-world scenarios, the process of gathering observations, validating hypotheses, and refining hypotheses is iterative. These simplified, static settings fail to adequately assess an agent’s capabilities in real, dynamic environments. Even in studies with interactive environments (Xu et al., 2024; Montes et al., 2022), information gathering, rule generation, and application are treated as independent objectives or are incomplete. In reality, these processes are dynamic and interdependent: generated hypotheses guide actions, and new observations refine those hypotheses. In contrast, RULEARN provides a realistic setting that allows LLMs to engage in dynamic, interdependent rule-learning processes, closely mirroring real-world scenarios. To effectively navigate and learn within this environment, IDEA equips LLM agents with the ability to emulate human reasoning by managing these interdependent processes.
@@ 3 The RULEARN Benchmark
We develop three puzzle sets, the Function Operator, the Escape Room, and the Reactor, each comprising 20 puzzles associated with environments of varying complexity.
The Function Operator. This puzzle type simulates scenarios where systemic theories or established knowledge (in this case, mathematics) is applicable. The agent interacts with a set of univariate multi-term equations involving integer parameters from [0,9] and elementary functions of the variable $x$, selected from $f(x)\in\{x^{0},x^{1},x^{2},\sin(x),\frac{1}{x},|x|,-x\}$. The agent is provided with the number of functions, the presence of specific parameters in each function (the exact numerical values of these parameters are unknown and represented by letters), and the types of elementary functions involved in the current puzzle. The goal of the agent is to deduce the values of these parameters. For example, in one puzzle, the ground truth is $\mathbf{F_{1}}(x)=a\sin(x)+b\times\frac{1}{x}$, $\mathbf{F_{2}}(x)=ax^{2}$ where $a=3$ and $b=2$. The agent knows the following information: There are three elementary functions in this puzzle $\{\sin(x),\frac{1}{x},x^{2}\}$, there are two functions $\mathbf{F_{1}}(x)$ and $\mathbf{F_{2}}(x)$, $\mathbf{F_{1}}(x)$ has 2 terms and parameters $a,b$ in it, and $\mathbf{F_{2}}(x)$ has 1 term and one parameter $a$. To solve the puzzle, the agent must interact with the environment through a defined action space: selecting a function and assigning values to $x$, then observing the resultant output. For example, assigning values 1 and 2 to $\mathbf{F_{2}}$ reveals a quadratic increase in output, indicating the presence of $x^{2}$ in $\mathbf{F_{2}}$. Similarly, assigning a value of 1 to $\mathbf{F_{1}}$ results in a floating-point output, rather than an integer, suggesting the inclusion of trigonometric components, confirming that $\sin(x)$ is a component of $\mathbf{F_{1}}$. Each puzzle features a different number of functions, unknown parameters, and elementary functions in use.
The Escape Room. This environment simulates scenarios where no established knowledge is applicable. We create a fictitious environment: an agent is in an escape room–an art gallery, tasked with deciphering a password to unlock a code-secured door and escape. The password is a 3-digit number where each digit represents the count of paintings of a specific type—watercolor, oil, or acrylic—that share a given color. The agent can explore the environment and investigate paintings, receiving brief descriptions such as This is an acrylic painting of a green jungle, providing information about the painting’s type and color. The agent only knows that the password is a 3-digit number and receives a hint to focus on a specific color relevant to the password. After forming an initial hypothesis and inputting the password into the door, the door provides hints indicating which digits are correct. Based on these hints, the agent refines its hypotheses and infers the relationship between the painting attributes and the passcode. To prevent the problem from being solved merely by brute force, the specific color associated with the password changes after every three failed attempts. Each puzzle presents a varying number of paintings; paintings in the same room as the agent are directly visible, while those in different rooms are not, requiring the agent to take actions such as moving to another gallery to reveal all the paintings.
The Reactor. This environment type also simulates scenarios where no established knowledge is applicable. However, unlike in the Escape Room where agents only input passwords to test hypotheses, in the Reactor, agents are challenged to design more fine-grained experiments to uncover hidden rules. Specifically, the agent’s task is to synthesize target strings using a reactor governed by a hidden string-combining rule. These strings are represented by sequences of alphabetic letters, such as A, B, AABB, and CAB. The reactor permits the agent to input two strings, initiating a reaction that produces a new string for use in subsequent experiments. The agent’s objective is to decipher the specific rules that govern string synthesis by methodically testing different string combinations, with the ultimate goal of synthesizing the target string using the discovered rules. We have designed four types of rules, detailed in Table 1.
The detailed statistics for each puzzle type and example puzzles are provided in Appendix A.2 and A.5, respectively. Note that the RULEARN benchmark is designed to simulate real-world interactions and is adaptable to complex text environments with diverse rules. To achieve this, we do not restrict the representation of rules to a specific formal language, as doing so would limit realism and complexity. Moreover, predefining the rule format would provide agents with prior knowledge, contradicting real-world conditions. Therefore, we enable the LLM agents to use natural language to describe rules in RULEARN, promoting generalizability and future extensions of the benchmark.

### v5

@@ Abstract
While large language models (LLMs) have been thoroughly evaluated for deductive and inductive reasoning, their proficiency in holistic rule learning in interactive environments remains less explored. We introduce RULEARN, a novel benchmark to assess the rule-learning abilities of LLM agents in interactive settings. In RULEARN, agents strategically interact with simulated environments to gather observations, discern patterns, and solve complex problems. To enhance the rule-learning capabilities for LLM agents, we propose IDEA, a novel reasoning framework that integrates the process of Induction, DEduction, and Abduction. The IDEA agent generates initial hypotheses from limited observations through abduction, devises plans to validate these hypotheses or leverages them to solve problems via deduction, and refines previous hypotheses through induction, dynamically establishing and applying rules that mimic human rule-learning behaviors. Our evaluation of the IDEA framework, which involves five representative LLMs, demonstrates significant improvements over the baseline. Furthermore, our study with human participants reveals notable discrepancies in rule-learning behaviors between humans and LLMs. We believe our benchmark will serve as a valuable and challenging resource, and IDEA will provide crucial insights for the development of LLM agents capable of human-like rule learning in real-world scenarios. Our code and data is publicly available [^1] https://github.com/KaiyuHe998/RULEARN_IDEA.
IDEA: Enhancing the Rule Learning Ability of Large Language Model Agent through Induction, Deduction, and Abduction
Kaiyu He, Mian Zhang, Shuo Yan, Peilin Wu, Zhiyu Zoey Chen Department of Computer Science University of Texas at Dallas {kaiyu.he, zhiyu.chen2}@utdallas.edu
@@ 1 Introduction
One major pillar of human intelligence is the ability to discern rules and apply them. We identify patterns, formulate hypotheses, and refine them by interacting with the environment. This exploratory process traditionally involves three stages: abduction, deduction, and induction. According to Charles Peirce’s definition (Frankfurt, 1958; Peirce, 1974), the rule-learning loop typically begins with an explanatory hypothesis that arises from abduction. This is followed by iterative experiments guided by the hypothesis, known as deduction, which leads to the modification and refinement of the hypothesis through induction (As shown in Figure 1). This establishes the foundation of how humans learn the rules of the world and apply them to effect change. In the real world, such a process involves iteratively refining hypotheses through interaction and evidence gathering, driven by a dynamic interplay between reasoning stages and the environment. However, recent studies often isolate abduction, induction, and deduction in non-interactive settings, missing this interdependent nature (Bowen et al., 2024; Wang et al., 2023; Saparov et al., 2024; Liu et al., 2024).
@@
Figure 2: A simplified puzzle in the RULEARN benchmark and the IDEA agent’s workflow (in real puzzles, agents have fewer initial observations and more complex rules). The agent generates an initial hypothesis through abduction, develops an exploration plan via deduction, and refines its hypothesis using induction. For example, the IDEA agent first hypothesizes that the password is the number of the blue paintings, tests this by entering code 003, and adjusts its strategy based on the feedback.
To address this gap, we introduce RULEARN, a new benchmark to evaluate the rule-learning abilities of LLM agents in interactive environments. RULEARN consists of 300 high-quality puzzles with hidden rules set in a text-based environment, where agents begin exploration without prior knowledge of these rules. Successfully solving the puzzles requires the agent to strategically select actions, efficiently gather pattern-revealing observations, and accurately reason from them to infer the hidden rules. RULEARN presents substantial challenges, as agents must rely entirely on observations generated by their chosen actions to discern rules. If these observations fail to reveal a clear pattern, the agent is likely to fail.
@@
We design three types of environments within RULEARN to evaluate the rule-learning ability in different scenarios: (1) The Function Operator: Determining the coefficients of mathematical functions defined by hidden expressions. Agents can assign various values to the input variables and observe the outputs, using this information to hypothesize the function’s form. The challenge lies in efficiently selecting input values that reveal the underlying structure and accurately computing the coefficients based on limited observations. (2) The Escape Room: Deciphering the passcode to exit an escape room. A hidden rule determines how the objects in the room infer the passcode digits. Agents interact with these objects to gather clues and input passcodes into the door. Based on feedback, agents formulate hypotheses and infer the relationship between the objects and the passcode. (3) The Reactor: Synthesizing target strings using a reactor with a hidden string-combining rule. Agents need to experiment with different inputs and analyze outputs to deduce the reactor’s transformation rule and achieve the desired outcome.
@@
To tackle the challenge in RULEARN, we introduce IDEA, a novel reasoning framework that integrates the process of Induction, DEduction, and Abduction. The IDEA agent employs these reasoning processes iteratively to explore the environments, learn rules, and achieve goals. In the abduction phase, the IDEA agent generates an initial hypothesis from limited observations. During the deduction phase, the IDEA agent creates and executes plans to attempt objectives or test its hypothesis. In the induction phase, the IDEA agent refines its hypothesis based on new observations, enhancing their accuracy and robustness. This iterative cycle enables the LLM agent to continually improve the learned rules through environmental feedback. An overview of how the IDEA agent solves puzzles in RULEARN is shown in Figure 2.
We evaluate IDEA on five popular LLMs—GPT-3.5-Turbo, GPT-4o, Gemma-7B, Llama3-8B, and Llama3-70B—observing roughly a 10% improvement in success rates compared to the baseline. Without hypothesis guidance, the baseline agent tends to choose direct, uninformed actions that fail to uncover the hidden rules. In contrast, the IDEA agent reduces repeated actions by 30.2%, obtains more diverse observations, and better understands the underlying rules. To further investigate their rule-learning capabilities, we compare LLM performance with that of 50 human participants. Although IDEA narrows the gap, LLMs still face challenges: (1) inefficient exploration in unfamiliar environments, resulting in insufficient evidence to reveal rules; (2) difficulty in deducing valid plans to verify current hypotheses and guide future exploration; and (3) reluctance to correct initial hypotheses when confronted with contradictory observations. These findings provide important insights into improving LLM agents to achieve more human-like rule learning in complex, real-world scenarios.
@@ 2 Related Works
Agents powered by large language models (LLMs) with memory, tool use, planning, and reasoning have shown notable progress in understanding complex tasks (Wang et al., 2024b; Chen et al., 2023; Zhou et al., 2023; Wang et al., 2024a; Andreas, 2022; Park et al., 2023; Zhong et al., 2023; Zhang et al., 2024b; Nakano et al., 2022; Lu et al., 2023; Shi et al., 2023a; Schick et al., 2023; Yuan et al., 2023; Shen et al., 2023; Yao et al., 2023a; Besta et al., 2024). Recent work examines different reasoning processes (abduction, deduction, induction) in LLMs (Bowen et al., 2024; Wang et al., 2023; Saparov et al., 2024; Cheng et al., 2024; Yang et al., 2024), but typically in isolation. As a result, their comprehensive rule-learning abilities in interactive settings remain underexplored.
Current reasoning tasks are hindered by inadequate benchmarks, primarily relying on QA datasets like Hotpot-QA (Yang et al., 2018) and Trivia-QA (Joshi et al., 2017) that lack the requirement for active information gathering, and coarse-grained interactive environments such as TextWorld (Côté et al., 2019) and AlfWorld (Shridhar et al., 2021) that limit agents to high-level actions (e.g., go to, open), impeding complex, experiment-driven rule discovery. This approach contrasts with real-world rule learning, which necessitates active evidence gathering, experimentation, and iterative refinement. Exacerbating this issue, many existing studies employ static, non-interactive scenarios where LLMs receive all information upfront (Yang et al., 2023; Zhu et al., 2024; Shi et al., 2023b; Liu et al., 2024), failing to mirror the dynamic nature of real-world learning processes. Furthermore, while recent efforts (Xu et al., 2024; Montes et al., 2022) attempt to integrate interactivity, they still treat information gathering, rule generation, and application as distinct phases, thereby undermining the development of agents capable of seamless integration of these elements in real-world settings.
This gap arises due to the absence of an interactive, text-based environment that supports fine-grained actions for rule learning. To address this, RULEARN enables actions where agents can input arbitrary strings, and our system parses these inputs, providing tailored feedback on a character-by-character basis. This capability allows LLM agents to conduct fine-grained experiments and closely mirror real-world rule learning. In turn, IDEA equips agents to effectively manage the interdependent processes of information gathering, hypothesis generation, and validation in a unified manner, offering a more comprehensive framework that mimics human rule-learning behaviors.
@@ 3 The RULEARN Benchmark
We develop three puzzle sets—Function Operator, Escape Room, and Reactor—each consisting of 100 unique puzzles of varying complexity, all manually created by our coauthors, who are computer science graduate students. Unlike existing fine-grained interactive environments, which are predominantly found in the robotics domain and offer significantly fewer tasks (Jain et al., 2020; Nasiriany et al., 2024; Zhang et al., 2024a), RULEARN is the first text-based environment providing such fine-grained interactions specifically for language agents.
@@
The Function Operator. This puzzle type simulates scenarios where systemic theories or established knowledge (in this case, mathematics) is applicable. The agent interacts with a set of univariate multi-term equations involving integer parameters from [0,9] and elementary functions of the variable $x$, selected from $f(x)\in\{x^{0},x^{1},x^{2},\sin(x),\frac{1}{x},|x|,-x\}$. The agent is provided with the number of functions, the presence of specific parameters in each function (the exact numerical values of these parameters are unknown and represented by letters), and the types of elementary functions involved in the current puzzle. The goal of the agent is to deduce the values of these parameters. For example, in one puzzle, the ground truth is $\mathbf{F_{1}}(x)=a\sin(x)+b\times\frac{1}{x}$, $\mathbf{F_{2}}(x)=ax^{2}$ where $a=3$ and $b=2$. The agent knows the following information: There are three elementary functions in this puzzle $\{\sin(x),\frac{1}{x},x^{2}\}$, there are two functions $\mathbf{F_{1}}(x)$ and $\mathbf{F_{2}}(x)$, $\mathbf{F_{1}}(x)$ has 2 terms and parameters $a,b$ in it, and $\mathbf{F_{2}}(x)$ has 1 term and one parameter $a$. To solve the puzzle, the agent must interact with the environment through a defined action space: selecting a function and assigning values to $x$, then observing the resultant output. For example, assigning values 1 and 2 to $\mathbf{F_{2}}$ reveals a quadratic increase in output, indicating the presence of $x^{2}$ in $\mathbf{F_{2}}$. Similarly, assigning a value of 1 to $\mathbf{F_{1}}$ results in a floating-point output, rather than an integer, suggesting the inclusion of trigonometric components, confirming that $\sin(x)$ is a component of $\mathbf{F_{1}}$. The difficulty of each puzzle is controlled by variations in the number of functions, unknown parameters, and elementary functions in use. We manually enumerated 100 combinations of functions, incorporating different numbers of terms and types of elementary functions to ensure a diverse range of puzzle complexity (see detailed distribution in Table 4 in Appendix  A.2).
The Escape Room. This environment simulates scenarios where no established knowledge is applicable. We create a fictitious setting: an agent is placed in an art gallery escape room and must decipher a 3-digit password to unlock a code-secured door. Each digit of the password represents the count of paintings of a specific type—watercolor, oil, or acrylic—that share a given color. The agent receives brief descriptions of paintings, such as “This is an acrylic painting of a green jungle”, indicating their type and color. Initially, the agent only knows the password is a 3-digit number and is given a hint about which color to focus on. After proposing a hypothesis and entering a password guess, the door provides feedback on which digits are correct, allowing the agent to refine its hypothesis. To prevent a brute-force approach, the specific color associated with the password changes after every three failed attempts. Each puzzle varies in the number of paintings, and while paintings in the same room are visible, those in other rooms remain hidden until the agent moves to access them. The difficulty of this puzzle type is controlled by the different number of paintings, whether agent need to as detailed in Table 5 in the Appendix  A.2.
The Reactor. This environment type also simulates scenarios where no established knowledge is applicable. However, unlike in the Escape Room where agents only input passwords to test hypotheses, in the Reactor, agents are challenged to design more fine-grained experiments to uncover hidden rules. Specifically, the agent’s task is to synthesize target strings using a reactor governed by a hidden string-combining rule. These strings are represented by sequences of alphabetic letters, such as A, B, AABB, and CAB. The reactor permits the agent to input two strings, initiating a reaction that produces a new string for use in subsequent experiments. The agent’s objective is to decipher the specific rules that govern string synthesis by methodically testing different string combinations, with the ultimate goal of synthesizing the target string using the discovered rules. We have designed four types of rules, detailed in Table 1. The difficulty of this puzzle type is controlled by the specific rules used, the length of the target string to be synthesized and the number of unique letters contained in the target string, as detailed in Table 6 in Appendix A.2.
The detailed statistics for each puzzle type and example puzzles are provided in Appendix A.2 and A.7, respectively. Note that the RULEARN benchmark is designed to simulate real-world interactions and is adaptable to complex text environments with diverse rules. To achieve this, we do not restrict the representation of rules to a specific formal language, as doing so would limit realism and complexity. Moreover, predefining the rule format would provide agents with prior knowledge, contradicting real-world conditions. Therefore, we enable the LLM agents to use natural language to describe rules in RULEARN, promoting generalizability and future extensions of the benchmark.